The amsintx package

American Mathematical Society Michael Downes

Version 0.9, 1995/01/27

\NeedsTeXFormat{LaTeX2e}
\errmessage{Dummy file amsintx.sty, release delayed}

The usual \endinput to ensure that random garbage at the end of the file doesn’t get copied by docstrip.

\endinput